# ResSVD: Residual Compensated SVD for Large Language Model Compression

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in a wide range of downstream natural language processing tasks. Nevertheless, their considerable sizes and memory demands hinder practical deployment, underscoring the importance of developing efficient compression strategies. Singular value decomposition (SVD) decomposes a matrix into orthogonal components, enabling efficient low-rank approximation. This is particularly suitable for LLM compression, where weight matrices often exhibit significant redundancy. However, current SVD-based methods neglect the residual matrix from truncation, resulting in significant truncation loss. Additionally, compressing all layers of the model results in severe performance degradation. To overcome these limitations, we propose ResSVD, a new post-training SVD-based LLM compression method. Specifically, we leverage the residual matrix generated during the truncation process to reduce truncation loss. Moreover, under a fixed overall compression ratio, we selectively compress the last few layers of the model, which mitigates error propagation and significantly improves the performance of compressed models. Comprehensive evaluations of ResSVD on diverse LLM families and multiple benchmark datasets indicate that ResSVD consistently achieves superior performance over existing counterpart methods, demonstrating its practical effectiveness.
@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools, delivering state-of-the-art performance across a wide range of tasks such as text generation, translation, and reasoning. The scaling law [18] has driven a trend toward increasingly large models, exemplified by models such as GPT [4], PaLM [6], LLaMA [38], Qwen [2], and Deepseek [25], which often contain tens to hundreds of billions of parameters. Despite their powerful capabilities, the enormous scale of LLMs poses serious challenges for efficient deployment due to high computational demands [48, 40, 35]. This resource burden not only limits deployment on edge devices and consumer-level hardware but also increases the cost and carbon footprint of serving LLMs in production [36, 31].
@@
As the scale of LLMs continues to grow, compression techniques, including weight quantization [11, 24, 16, 23], network pruning [27, 10, 37, 12], knowledge distillation [13, 43, 42, 47], and low-rank decomposition [15, 19, 44, 41, 22], have become increasingly important for the practical deployment of LLMs in resource-constrained environments. The syntactic and semantic correlations acquired during training often induce redundancy in LLM weight matrices, giving rise to a low-rank structure [33]. As a result, singular value decomposition (SVD) provides a principled and effective approach for compressing these matrices with minimal loss. In particular, post-training approaches are gaining traction, as they can significantly reduce memory and compute requirements without the need for expensive retraining, making them especially suitable for scaling up foundation models. Recent studies on post-training SVD-based LLM compression, including ASVD [44], SVD-LLM [41], and AdaSVD [22], have made significant progress in reducing model size while preserving performance, demonstrating the effectiveness of low-rank approximation techniques in compressing LLMs. More precisely, while both ASVD and SVD-LLM apply weight matrix scaling before SVD truncation, SVD-LLM outperforms ASVD by leveraging a data whitening technique that enables a direct mapping between singular values and truncation loss. As a further advancement, AdaSVD compensates for truncation loss by iteratively updating singular matrices. However, existing methods suffer from two major limitations. First, existing methods ignore the importance of the residual matrix generated during SVD truncation, leading to significant truncation loss, as indicated by our theoretical analysis in Section 3.1. Second, compressing all model layers often results in high layer-wise error and severe error propagation, as confirmed by our results in Figure 3.
@@
In this work, we propose ResSVD, a new post-training SVD-based compression method for LLMs. Building upon the two key observations outlined above, ResSVD introduces two core technical innovations. ① Residual compensation for SVD truncation: The residual matrix produced during SVD truncation can be effectively utilized to reduce the truncation loss. Specifically, we perform SVD truncation in two stages: we first truncate the original weight matrix $\boldsymbol{W}$ to obtain its low-rank approximation $\boldsymbol{W}_{r_{1}}$. After that, we compute the residual matrix $\boldsymbol{R}$ between $\boldsymbol{W}$ and $\boldsymbol{W}_{r_{1}}$. Second, we apply SVD truncation to $\boldsymbol{R}$, yielding $\boldsymbol{R}_{r_{2}}$. Finally, we construct the compressed weight matrix $\boldsymbol{\hat{W}}_{r}=\boldsymbol{W}_{r_{1}}+\boldsymbol{R}_{r_{2}}$, where the total rank satisfies $r_{1}+r_{2}=r$. Theoretical justification is provided in Section 3.1. ② Partial-layer compression for SVD: LLMs consist of a sequence of consecutive layers, where the output of each layer serves as the input to the next. As a result, any error introduced in earlier layers can propagate and accumulate through subsequent layers, leading to severe degradation in performance. To mitigate this, we propose compressing only the last few layers under a fixed overall compression ratio while keeping the earlier layers intact. This strategy ensures that the earlier layers remain error-free, thereby reducing the impact of error propagation.
Our key contributions can be summarized as follows:
@@
  • Extensive experiments on multiple LLMs (LLaMA, OPT, Mistral, and Vicuna) and benchmark datasets (both language modeling and zero-shot reasoning) demonstrate that ResSVD outperforms existing methods across a wide range of compression ratios.
@@ 2.1 Techniques for Large Language Model Compression
The growing size of large language models (LLMs) has raised increasing concerns over their computational and memory demands. To address these challenges, a variety of model compression techniques have been proposed. Conventional approaches often require computationally expensive retraining, which is generally impractical due to the substantial computational cost associated with the massive size of LLMs. Consequently, recent efforts have shifted toward more resource-friendly post-training compression techniques [10, 11, 49]. Typically employed approaches include network pruning and weight quantization. And pruning techniques can be classified into unstructured and structured methods. Unstructured pruning removes individual weights based on importance scores. For example, SparseGPT [10] performs one-shot pruning using second-order approximations without retraining. However, since unstructured pruning retains the original matrix shape, it offers limited inference acceleration and requires specialized hardware. In contrast, structured pruning eliminates entire blocks or channels, enabling compatibility with conventional hardware platforms. LLM-Pruner [27] groups dependent linear projections into coupled structures, assigns each group a loss-aware importance score, prunes the least important groups, and applies LoRA fine-tuning to restore performance. Additionally, ZipLM [20] prioritizes pruning components that yield the worst trade-off between latency and loss, but this often causes notable performance degradation. Quantization provides another mainstream solution. GPTQ [11] gradually quantizes and updates each weight using the Hessian matrix to minimize the quantization error. AWQ [24] preserves important weight channels by selecting reparameterization coefficients via grid search. Furthermore, BiLLM [16] and ARB-LLM [23] push quantization to the 1-bit level, while still delivering impressive accuracy across downstream tasks. However, these techniques still cause significant performance degradation, especially at low bit-widths, due to the lack of weight and activation adaptation.
@@ 2.2 SVD-based Techniques for Compressing LLMs
Singular value decomposition (SVD) is a commonly employed technique for reducing matrix dimensionality by representing the original matrix as the product of two low-rank factor matrices. Recent research has demonstrated the potential of SVD-based LLM compression methods, but comprehensive exploration remains limited. The conventional SVD only aims to compress the original weight matrix, without accounting for the relative importance of individual parameters, leading to significant approximation errors. To address this issue, Hsu et al. [15] propose FWSVD, incorporating Fisher information to weight the importance of parameters. However, it relies on computationally intensive training and was originally applied only to small language models (e.g., BERT [9], ALBERT [21]). LoRD [19] first applies SVD to LLMs by grouping layers to improve efficiency. However, similar to conventional SVD, LoRD shares the limitation of overlooking input activations. To address this problem, Yuan et al. [44] introduce ASVD, which mitigates the impact of activation outliers by reshaping the weight matrix based on activation distribution, thereby enhancing the precision of the low-rank decomposition. Additionally, SVD-LLM [41] establishes a direct mapping between singular values and truncation loss, which means that truncating the smallest singular values leads to minimal truncation loss using a truncation-aware data whitening technique. More recently, AdaSVD [22] proposes an adaptive compensation method, iteratively updating matrices during truncation. Despite these advances, all existing methods ignore the residual matrix produced during SVD truncation, which can significantly compensate for the SVD truncation loss. Moreover, compressing only the last few layers of the model under a fixed target compression ratio can provide better performance for compressed models. However, existing methods compress all layers, either apply a uniform compression ratio across all layers or assign variable ratios based on layer-wise importance, which often results in sub-optimal performance.
3 ResSVD
@@
Figure 1: Framework of ResSVD. The last $k$ layers are selected through partial-layer compression and compressed using residual compensation with calibration data. The first $(N-k)$ layers remain intact, while the last $k$ layers are replaced by low-rank approximations.
@@
The framework of ResSVD is illustrated in Figure 1. We first perform SVD to compress the last few layers of the model, while ensuring that the overall compression ratio is satisfied, and compute the corresponding residual matrix. A second SVD is then applied to the residual matrix to obtain its low-rank approximation. The two truncated matrices are subsequently combined to construct the final compressed weight matrices. In Section 3.1, we first provide a theoretical analysis demonstrating how residual compensation effectively reduces truncation loss. Subsequently, in Section 3.2, we explain the benefits of partial-layer compression in mitigating error propagation. The pseudocode of ResSVD is shown in Algorithm 1, and the pseudocodes for residual compensation and partial-layer compression are provided in Appendix A.1.
Algorithm 1 Pseudocode of ResSVD
@@
0:  Original LLM: $M$, weight matrix: $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, number of model layers: $N$, intermediate rank $r_{1}$, step: $s$
0:  Compressed LLM $M^{\prime}$ by ResSVD
1:  $CD\leftarrow$ Randomly collect calibration samples from the dataset
2:  $\mathrm{Set}_{\boldsymbol{W}}\leftarrow M$, $\mathrm{Set}_{\boldsymbol{W^{\prime}}}\leftarrow\emptyset$triangleright Initialize the sets of weight matrices
3:  $k$, $R_{l}\leftarrow$ Partial-layer Compression($M$, $N$, $R_{o}$, $s$)
@@
6:        $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)\leftarrow\mathrm{Set}_{\boldsymbol{W}}(i)$triangleright Current weight matrices stay the same
@@
8:        $\boldsymbol{W}_{i}\leftarrow$ $\mathrm{Set}_{\boldsymbol{W}}(i)$, ${\boldsymbol{S}_{i}}\leftarrow$ Whitening(${\boldsymbol{W}_{i}}$, $CD$) $\triangleright$ Initialize weight matrices to compress
9:        $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)\leftarrow$ Residual Compensation($S_{i}$, $k$, $R_{l}$, $r_{1}$)
@@
11:     $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$leftarrow $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)$cup $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$triangleright Append weight matrices after operation
@@
13:  $M^{\prime}$leftarrow Update($M$, $CD$, $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$)
@@
3.1 Residual Compensation for SVD Truncation
Preliminaries for SVD. Typical SVD-based LLM compression methods apply SVD to the original weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, and discard the smallest singular values to obtain a compressed low-rank approximation $\boldsymbol{W}_{r}$:
where $\boldsymbol{U}_{r}\in\mathbb{R}^{m\times r}$ and $\boldsymbol{V}_{r}^{T}\in\mathbb{R}^{r\times n}$ are composed of the top-$r$ left and right singular vectors, respectively, and $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the corresponding singular values. Following prior post-training SVD-based works [44, 41], the optimization objective for SVD truncation in LLMs can be formulated as:
where $\boldsymbol{X}$ denotes the activation of $\boldsymbol{W}$ given an input, and $\mathcal{L}$ is the truncation loss measured by the Frobenius norm. Although previous works [15, 44, 41] have made significant progress in minimizing $\mathcal{L}$, they consistently overlook the residual matrix generated during truncation, despite its potential to compensate for the loss introduced by low-rank approximation.
Based on Equation 3, minimizing the truncation loss reduces to minimizing the discrepancy between the original weight matrix $\boldsymbol{W}$ and its low-rank approximation $\boldsymbol{W}_{r}$. Accordingly, we reformulate the optimization objective as:
@@ 3.1 Residual Compensation for SVD Truncation
Method Explanation for ResSVD. Given the original weight matrix $\boldsymbol{W}$ and the target rank $r$, we decompose $r$ into two components: an intermediate rank $r_{1}$ and a residual rank $r_{2}=r-r_{1}$. The compression process consists of two SVD truncation steps. First, we apply SVD to $\boldsymbol{W}$ and retain top-$r_{1}$ singular values to obtain a low-rank approximation: $\boldsymbol{W}_{r_{1}}=\boldsymbol{U}_{r_{1}}\boldsymbol{\Sigma}_{r_{1}}\boldsymbol{V}_{r_{1}}^{T}$. The residual matrix is then computed as the difference between the original matrix and its rank-$r_{1}$ approximation: $\boldsymbol{R}=\boldsymbol{W}-\boldsymbol{W}_{r_{1}}$. We perform a second SVD on $\boldsymbol{R}$ and retain the top-$r_{2}$ singular values: $\boldsymbol{R}_{r_{2}}=\boldsymbol{U}_{r_{2}}\boldsymbol{\Sigma}_{r_{2}}\boldsymbol{V}_{r_{2}}^{T}$. Finally, we obtain the compressed weight matrix by combining the two approximations: $\hat{\boldsymbol{W}}_{r}=\boldsymbol{W}_{r_{1}}+\boldsymbol{R}_{r_{2}}=\boldsymbol{U}_{r}\boldsymbol{\Sigma}_{r}\boldsymbol{V}_{r}^{T}$.
@@
In this way, the residual-compensated compressed matrix $\hat{\boldsymbol{W}}_{r}$ provides a closer approximation to the original weight matrix $\boldsymbol{W}$ than the directly truncated matrix $\boldsymbol{W}_{r}$, as proven in the theoretical analysis that follows.
Lemma 3.1.
@@
Eckart-Young-Mirsky Theorem [14]. Given a matrix $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ with singular values $\sigma_{1}(\boldsymbol{A})\geq\sigma_{2}(\boldsymbol{A})\geq\sigma_{3}(\boldsymbol{A})\geq...\geq\sigma_{\min(m,n)}(\boldsymbol{A})$. The best rank-$r$ approximation $\boldsymbol{A}_{r}$ is given by:
$\boldsymbol{A}_{r}=\sum_{i=1}^{r}\sigma_{i}(\boldsymbol{A})u_{i}(\boldsymbol{A})v_{i}^{T}(\boldsymbol{A}),$ (5)
and the truncation loss satisfies:
$\big{\|}\boldsymbol{A}-\boldsymbol{A}_{r}\big{\|}_{F}^{2}=\sum_{i=r+1}^{\min(m,n)}\sigma_{i}^{2}(\boldsymbol{A}).$ (6)
As defined above, $\hat{\boldsymbol{W}}_{r}=\boldsymbol{W}_{r_{1}}+\boldsymbol{R}_{r_{2}}$. Let $\mathcal{L}_{C}$ denote the truncation loss between the original matrix $\boldsymbol{W}$ and the residual-compensated compressed matrix $\hat{\boldsymbol{W}}_{r}$. Then, by applying Lemma 3.1, we have:
$\mathcal{L}_{C}=\big{\|}\boldsymbol{W}-\hat{\boldsymbol{W}}_{r}\big{\|}_{F}^{2}=\big{\|}\boldsymbol{W}-(\boldsymbol{W}_{r_{1}}+\boldsymbol{R}_{r_{2}})\big{\|}_{F}^{2}=\big{\|}\boldsymbol{R}-\boldsymbol{R}_{r_{2}}\big{\|}_{F}^{2}=\sum_{i=r_{2}+1}^{rank(\boldsymbol{R})}\sigma_{i}^{2}(\boldsymbol{R}),$ (7)
$rank(\boldsymbol{R})\leq\min(m,n)-r_{1}.$
Lemma 3.2.
If $\boldsymbol{A}_{r}$ is the best rank-r approximation of $\boldsymbol{A}\in\mathbb{R}^{m\times n}$, and the residual is $\boldsymbol{B}=\boldsymbol{A}-\boldsymbol{A}_{r}$. The we have:
$\sigma_{i}(\boldsymbol{B})\leq\sigma_{r+i}(\boldsymbol{A}),1\leq i\leq\min(m,n)-r.$ (8)
@@
Theorem 3.3. If $\boldsymbol{W}_{r}$ is the rank-$r$ direct truncation matrix form the original matrix $\boldsymbol{W}$, $\hat{\boldsymbol{W}}_{r}$ is the residual-compensated compressed matrix, then $\hat{\boldsymbol{W}}_{r}$ is closer to $\boldsymbol{W}$.

### v3

# ERC-SVD: Error-Controlled SVD for Large Language Model Compression

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in a wide range of downstream natural language processing tasks. Nevertheless, their considerable sizes and memory demands hinder practical deployment, underscoring the importance of developing efficient compression strategies. Singular value decomposition (SVD) decomposes a matrix into orthogonal components, enabling efficient low-rank approximation. This is particularly suitable for LLM compression, where weight matrices often exhibit significant redundancy. However, current SVD-based methods neglect the residual matrix from truncation, resulting in significant truncation loss. Additionally, compressing all layers of the model results in severe error propagation. To overcome these limitations, we propose ERC-SVD, a new post-training SVD-based LLM compression method from an error-controlled perspective. Specifically, we leverage the residual matrix generated during the truncation process to reduce truncation loss. Moreover, under a fixed overall compression ratio, we selectively compress the last few layers of the model, which mitigates error propagation and improves compressed model performance. Comprehensive evaluations on diverse LLM families and multiple benchmark datasets indicate that ERC-SVD consistently achieves superior performance over existing counterpart methods, demonstrating its practical effectiveness.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 1: Left: The Kendall correlation between the final-layer error and the average zero-shot accuracy. Middle: Layer-wise error comparison between ERC-SVD and SVD-LLM of LLaMA-2-7B under 20% compression ratio. Right: The accuracy comparison of LLaMA-2-7B compressed by different methods under 20% compression ratio on seven reasoning and understanding tasks.
@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools, delivering state-of-the-art performance across a wide range of tasks such as text generation, translation, and reasoning. The scaling law [kaplan2020scaling] has driven a trend toward increasingly large models, exemplified by models such as GPT [brown2020language], PaLM [chowdhery2023palm], LLaMA [touvron2023llama], Deepseek [liu2024deepseek], and Qwen [yang2025qwen3], which often contain tens to hundreds of billions of parameters. Despite their powerful capabilities, the enormous scale of LLMs poses serious challenges for efficient deployment due to high computational demands [sheng2023flexgen, zhou2024survey, wang2024model]. This resource burden not only limits deployment on edge devices and consumer-level hardware but also increases the cost and carbon footprint of serving LLMs in production [strubell2020energy, patterson2022carbon].
@@
As the scale of LLMs continues to grow, compression techniques, including weight quantization [frantar2022gptq, lin2024awq, huang2024billm, li2024arb], network pruning [ma2023llm, frantar2023sparsegpt, sunsimple, gao2024disp], knowledge distillation [gu2023minillm, yang2024survey, xu2024survey, zhang2024dual], and low-rank decomposition [hsulanguage, kaushal2023lord, yuan2023asvd, wang2024svd, li2025adasvd], have become increasingly important for the practical deployment of LLMs in resource-constrained environments. The syntactic and semantic correlations acquired during training often induce redundancy in LLM weight matrices, giving rise to a low-rank structure [saha2024compressing]. As a result, singular value decomposition (SVD) provides a principled and effective approach for compressing these matrices with minimal loss. In particular, post-training approaches are gaining traction, as they can significantly reduce memory and compute requirements without the need for expensive retraining, making them especially suitable for scaling up foundation models. Recent studies on post-training SVD-based LLM compression, including ASVD [yuan2023asvd], SVD-LLM [wang2024svd], Basis Sharing [wang2024basis], and AdaSVD [li2025adasvd], have made significant progress in reducing model size while preserving performance, demonstrating the effectiveness of low-rank approximation techniques in compressing LLMs. However, existing methods suffer from two major limitations. First, existing methods ignore the importance of the residual matrix generated during SVD truncation, leading to significant truncation loss. Second, compressing all model layers often results in high layer-wise error and severe error propagation.
@@
In this work, we propose ERC-SVD, a new post-training SVD-based compression method for LLMs. Building upon the key observations outlined above, ERC-SVD introduces two core technical innovations. ① Residual compensation for SVD truncation: The residual matrix produced during SVD truncation can be effectively utilised to reduce the truncation loss. Specifically, we perform SVD truncation in two stages: we first truncate the original weight matrix $\boldsymbol{W}$ to obtain its intermediate low-rank approximation $\boldsymbol{W}_{r_{i}}$. After that, we compute the residual matrix $\boldsymbol{R}$ between $\boldsymbol{W}$ and $\boldsymbol{W}_{r_{i}}$. Second, we apply SVD truncation to $\boldsymbol{R}$, yielding $\boldsymbol{R}_{r_{r}}$. Finally, we construct the compressed weight matrix $\boldsymbol{\hat{W}}_{r}=\boldsymbol{W}_{r_{i}}+\boldsymbol{R}_{r_{r}}$, where the rank satisfies $r_{i}+r_{r}=r$. Detailed description and a mathematical proof are provided in Section 3.1. ② Partial-layer compression for SVD: LLMs consist of a sequence of consecutive layers, where the output of each layer serves as the input to the next. Therefore, any error introduced in earlier layers can propagate and accumulate through subsequent layers, leading to severe performance degradation. To mitigate this, we propose compressing only the last few layers under a fixed overall compression ratio while keeping the earlier layers intact. This strategy ensures that the earlier layers remain error-free, reducing the impact of error propagation.
Our key contributions can be summarised as follows:
@@
  • Extensive experiments on multiple LLMs (LLaMA, OPT, Mistral, Vicuna, and Qwen) and benchmark datasets (both language modelling and zero-shot reasoning) demonstrate that ERC-SVD outperforms existing methods across a wide range of compression ratios.
@@ 2.1 Techniques for Large Language Model Compression
The growing size of large language models (LLMs) has raised increasing concerns over their computational and memory demands. To address these challenges, a variety of model compression techniques have been proposed. Conventional approaches often require computationally expensive retraining, which is generally impractical due to the substantial computational cost associated with the massive size of LLMs. Consequently, recent efforts have shifted toward more resource-friendly post-training compression techniques [frantar2022gptq, frantar2023sparsegpt, zhu2024survey]. Typically employed approaches include network pruning and weight quantization. And pruning techniques can be classified into unstructured and structured methods. Unstructured pruning removes individual weights based on importance scores. For example, SparseGPT [frantar2023sparsegpt] performs one-shot pruning using second-order approximations without retraining. However, since unstructured pruning retains the original matrix shape, it offers limited inference acceleration and requires specialized hardware. In contrast, structured pruning eliminates entire blocks or channels, enabling compatibility with conventional hardware platforms. LLM-Pruner [ma2023llm] groups dependent linear projections into coupled structures, assigns each group a loss‑aware importance score, prunes the least important groups, and applies LoRA fine‑tuning to restore performance. Additionally, ZipLM [kurtic2023ziplm] prioritizes pruning components that yield the worst trade-off between latency and loss, but this often causes notable performance degradation. Quantization provides another mainstream solution. GPTQ [frantar2022gptq] gradually quantizes and updates each weight using the Hessian matrix to minimize the quantization error. AWQ [lin2024awq] preserves important weight channels by selecting reparameterization coefficients via grid search. Furthermore, BiLLM [huang2024billm] and ARB-LLM [li2024arb] push quantization to the 1-bit level, while still delivering impressive downstream task accuracy. However, these techniques still cause significant performance degradation, especially at low bit-widths, due to the lack of weight and activation adaptation.
@@ 2.2 SVD-based Techniques for Compressing LLMs
Singular value decomposition (SVD) is a commonly employed technique for reducing matrix dimensionality by representing the original matrix as the product of two low-rank factor matrices. Recent research has demonstrated the potential of SVD-based LLM compression methods, yet comprehensive exploration remains limited. Hsu et al. [hsulanguage] propose FWSVD, incorporating Fisher information to weight the importance of parameters. However, it relies on computationally intensive training and was originally applied only to small language models (e.g., BERT [devlin2019bert], ALBERT [lan2019albert]). LoRD [kaushal2023lord] first applies SVD to LLMs by grouping layers to improve efficiency. However, LoRD overlooks the importance of input activations. To address this limitation, Yuan et al. [yuan2023asvd] introduce ASVD, which mitigates the impact of activation outliers by reshaping the weight matrix based on activation distribution, thereby enhancing the precision of the low-rank decomposition. CALDERA [saha2024compressing] exploits the inherent low-rank structure of weight matrices by approximating them through a low-rank, low-precision decomposition. Additionally, SVD-LLM [wang2024svd] establishes a direct mapping between singular values and truncation loss, which means that truncating the smallest singular values leads to minimal truncation loss. Basis Sharing [wang2024basis] achieves more effective LLM compression through sharing parameters across different layers. More recently, AdaSVD [li2025adasvd] proposes an adaptive compensation method, iteratively updating matrices during truncation.
Despite these advances, all existing methods ignore the residual matrix produced during SVD truncation, which can significantly compensate for the SVD truncation loss. Moreover, compressing only the last few layers of the model under a fixed target compression ratio can provide better performance for compressed models. However, existing methods compress all layers, either apply a uniform compression ratio across all layers or assign variable ratios based on layer-wise importance, which often results in sub-optimal performance.
3 ERC-SVD
@@
Figure 2: The overall framework of ERC-SVD, and comparison with other methods. The last $k$ layers are selected through partial-layer compression and compressed using residual compensation with calibration data. Refer to caption denotes these layers remain intact, while Refer to caption denotes these layers are replaced by low-rank approximations. The overall compression ratio is $R_{o}$, for ERC-SVD, the first (Nk)(N-k) layers stay unchanged, and the layer compression ratio $R_{l}$ for last $k$ layers is $(N\cdot R_{o})/k$.
@@
The framework of ERC-SVD is illustrated in Figure 2. We first perform SVD to compress the last few layers of the model, while ensuring that the overall compression ratio is satisfied, and compute the corresponding residual matrix. A second SVD is then applied to the residual matrix to obtain its low-rank approximation. The two truncated matrices are subsequently combined to construct the final compressed weight matrices. Compared with other methods [yuan2023asvd, wang2024svd, wang2024basis, li2025adasvd], we utilize the residual matrix during the truncation process to compensate for the truncation loss. In Section 3.1, we first provide a detailed description of how residual compensation effectively works. Subsequently, in Section 3.2, we explain the benefits of partial-layer compression in mitigating error propagation. The pseudocode of ERC-SVD is shown in Algorithm 1, with the pseudocode for residual compensation and partial-layer compression provided in Algorithm 2 and Algorithm 3, respectively.
3.1 Residual Compensation for SVD Truncation
Preliminaries for SVD.
SVD-based compression methods apply SVD to the original weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, discarding the smallest singular values to obtain its low-rank approximation $\boldsymbol{W}_{r}$:
where $\boldsymbol{U}_{r}\in\mathbb{R}^{m\times r}$ and $\boldsymbol{V}_{r}^{T}\in\mathbb{R}^{r\times n}$ are composed of the top-$r$ left and right singular vectors, respectively, and $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the corresponding singular values. Following prior post-training SVD-based works [yuan2023asvd, wang2024svd], the optimization objective for SVD truncation in LLMs can be formulated as:
where $\boldsymbol{X}$ denotes the activation of $\boldsymbol{W}$ given an input, and $\mathcal{L}$ is the truncation loss measured by the Frobenius norm. Although previous works [hsulanguage, yuan2023asvd, wang2024svd] have made significant progress in minimizing $\mathcal{L}$, they consistently overlook the residual matrix generated. Based on Equation 3, minimizing the truncation loss reduces to minimizing the discrepancy between the original weight matrix $\boldsymbol{W}$ and its low-rank approximation $\boldsymbol{W}_{r}$. Accordingly, we reformulate the optimization objective as:
Algorithm 1 Pseudocode of ERC-SVD
@@ 3.1 Residual Compensation for SVD Truncation
0: Original LLM: $M$, weight matrix: $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, number of model layers: NN, residual compensation rank $r_{r}$, step: $s$
0: Compressed LLM $M^{\prime}$ by ERC-SVD
[…]
3:$k$, $R_{l}\leftarrow$ Partial-layer Compression($M$, NN, $R_{o}$, $s$)
@@
6:   $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)\leftarrow\mathrm{Set}_{\boldsymbol{W}}(i)$ $\triangleright$ Current weight matrices stay the same
@@
8:   $\boldsymbol{W}_{i}\leftarrow$ $\mathrm{Set}_{\boldsymbol{W}}(i)$ $\triangleright$ Initialize weight matrices to compress
9:   $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)\leftarrow$ Residual Compensation($M$, $\boldsymbol{W}_{i}$, $R_{l}$, $r_{r}$)
@@
11:  $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$ $\leftarrow$ $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)$ $\cup$ $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$ $\triangleright$ Append weight matrices after operation
@@
[…]
@@
Explanation for Residual Compensation.
Given the original weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$ and the layer compression ratio $R_{l}$, the compression ratio in this work denotes the fraction of parameters removed. We define the scale of the matrix as: $\alpha=(m\cdot n)/(m+n)$, where $m$ and $n$ represent the input and output dimensions of the matrix, hence, the target rank for each layer is $r=(1-R_{l})\cdot\alpha$. We introduce a residual compensation factor $\beta$, a hyperparameter fixed at $0.05$ in all experiments. The target rank $r$ is decomposed into two components: a residual compensation rank $r_{r}$, defined as $r_{r}=\alpha\cdot\beta$, and an intermediate rank $r_{i}$, defined as $r_{i}=r-r_{r}$.
@@
The entire compression process comprises two SVD truncations. First, we apply SVD to $\boldsymbol{WS}$, where $\boldsymbol{S}$ is the scaling weight matrix obtained from the input activation, and retain the top-$r_{i}$ singular values: $\boldsymbol{WS}\approx\boldsymbol{W}_{r_{i}}^{\prime}=\boldsymbol{U}_{r_{i}}\boldsymbol{\Sigma}_{r_{i}}\boldsymbol{V}_{r_{i}}^{{}^{\prime}T}$, and multiply $\boldsymbol{S^{-1}}$ to obtain an intermediate low-rank approximation of the original weight matrix: $\boldsymbol{W}_{r_{i}}=\boldsymbol{W}_{r_{i}}^{\prime}\boldsymbol{S^{-1}}=\boldsymbol{U}_{r_{i}}\boldsymbol{\Sigma}_{r_{i}}\boldsymbol{V}_{r_{i}}^{T}.$ The residual matrix $\boldsymbol{R}$ is computed as the difference between $\boldsymbol{W}$ and $\boldsymbol{W}_{r_{i}}$, where $\boldsymbol{R}=\boldsymbol{W}-\boldsymbol{W}_{r_{i}}.$ We then perform a second SVD on $\boldsymbol{R}$ and retain its top-$r_{r}$ singular values: $\boldsymbol{R}_{r_{r}}=\boldsymbol{U}_{r_{r}}\boldsymbol{\Sigma}_{r_{r}}\boldsymbol{V}_{r_{r}}^{T}$. Finally, we construct the compressed weight matrix by combining the two approximations:
$\hat{\boldsymbol{W}}_{r}=\boldsymbol{W}_{r_{i}}+\boldsymbol{R}_{r_{r}}=\boldsymbol{U}_{r}\boldsymbol{\Sigma}_{r}\boldsymbol{V}_{r}^{T}=\hat{\boldsymbol{U}}_{r}\hat{\boldsymbol{V}}_{r}.$ (5)
@@
In this way, the residual-compensated compressed matrix $\hat{\boldsymbol{W}}_{r}$ provides a closer approximation to the original weight matrix $\boldsymbol{W}$ than the directly truncated matrix $\boldsymbol{W}_{r}$, as proven in the following.
Lemma 1.
@@
Eckart-Young-Mirsky Theorem [horn2012matrix]. Let $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ be a matrix and $r\leq\min(m,n)$ be a given rank. If $\boldsymbol{A}_{r}$ denotes the optimal rank-$r$ approximation of $\boldsymbol{A}$ by SVD, then for any matrix $\boldsymbol{B}$ of rank $r$, the following inequality holds, demonstrating that $\boldsymbol{A}_{r}$ is the optimal rank-$r$ approximation of $\boldsymbol{A}$:
$\big\|\boldsymbol{A}-\boldsymbol{A}_{r}\big\|_{F}^{2}\leq\big\|\boldsymbol{A}-\boldsymbol{B}\big\|_{F}^{2}.$ (6)
Lemma 2.
Let $\boldsymbol{W}^{\prime}$ be the rank-$r$ approximation defined by the weighted low-rank reconstruction $\boldsymbol{W}^{\prime}=SVD_{r}(\boldsymbol{WS})\boldsymbol{S}^{-1}$. A direct consequence of Lemma 1 is the following inequality:
$\big\|\boldsymbol{W}-SVD_{r}(\boldsymbol{W})\big\|_{F}^{2}\leq\big\|\boldsymbol{W}-\boldsymbol{W}^{\prime}\big\|_{F}^{2},$ (7)
@@
Theorem 3. If $\boldsymbol{W}_{r}$ is the rank-$r$ direct truncation of the original matrix $\boldsymbol{W}$, and $\hat{\boldsymbol{W}}_{r}$ is the residual-compensated compressed matrix, where $r=r_{i}+r_{r}$. Then, under the weight reconstruction error, ERC-SVD is superior to the direct truncation method.